OpenStudy (anonymous):

@TheSmartOne @iGreen. What is the difference : 4/3-7/9

OpenStudy (anonymous):

@Luigi0210

OpenStudy (anonymous):

?

OpenStudy (luigi0210):

Get common denominators: \(\Large \frac{4*3}{3*3}=\frac{12}{9}\) Now doo ett \(\Large \frac{12}{9}-\frac{7}{9}\)

OpenStudy (anonymous):

5/9?

TheSmartOne (thesmartone):

Correct.

TheSmartOne (thesmartone):

@zkmath @~

OpenStudy (anonymous):

new one okay?

TheSmartOne (thesmartone):

okay..